# Controlling Continuous Relaxation for Combinatorial Optimization

### v3

@@ 3 Continuous relaxation annealing for UL-based solvers
In this section, we first discuss the practical issues of UL-based solvers and then introduce continuous relaxation annealing (CRA) as a solution to these issues.
@@ 3.1 Motivation: practical issues of UL-based solvers
UL-based solvers, for a single instance CO problems, are effective for addressing large-scale CO problems. However, they present following two practical issues, indicated in several previous studies (Wang and Li, 2023; Angelini and Ricci-Tersenghi, 2023). In addition, we also numerically validate these issues; see Appendix F.1 for detailed results.
@@
UL-based solvers employ a continuous relaxation strategy to train NNs and then convert the relaxed continuous variables into discrete binary variables through artificial rounding as discussed in Section 2.1. This inherent ambiguity in the continuous relaxation strategy often leads to potential discrepancies between the optimal solutions of the original discrete CO problem and those of the relaxed continuous versions. The continuous relaxation expands the solution space, often leading to the emergence of continuous values that reduce the cost compared to an optimal binary value. Indeed, while linear relaxation can provide an optimal solution for original discrete problems on bipartite graphs (Hoffman and Kruskal, 2010), it typically leads to solutions with $\nicefrac{{1}}{{2}}$ values, which is known to half-integrality (Nemhauser and Trotter Jr, 1974). Existing rounding methods lose their robustness in these scenarios. In deed, PI-GNN solver often outputs values close to $\nicefrac{{1}}{{2}}$, demonstrating the limitations of current rounding techniques for UL-based solvers.
@@
Recently, Angelini and Ricci-Tersenghi (2023) demonstrated that PI-GNN solver could not achieve results comparable to those of the degree-based greedy algorithm (DGA) (Angelini and Ricci-Tersenghi, 2019) in solving the MIS problems on RRGs. Angelini and Ricci-Tersenghi (2023) also emphasized the importance of evaluating UL-based solvers in complex CO problems, where the performance of greedy algorithms worsens. The representative example is MIS problems on RRGs with a constant degree $d>16$, where a clustering transition occurs in the solution space, forming barriers that impede optimization. Moreover, Wang and Li (2023) emphasized the importance of using training/historical datasets, ${\mathcal{D}}=\{C_{\mu}\}_{1\leq\mu\leq P}$, which contains various graphs and initialization using the outputs of greedy solvers, e.g., DGA and RGA for MIS problems. Their numerical analysis indicated that PI-GNN solver tends to be trapped in local optima when directly optimized for a single instance without leveraging such a training dataset. However, in a practical setting, systematic methods for generating or collecting training datasets, ${\mathcal{D}}$, that effectively avoid local optima remains unclear. The extra training for instances that do not contribute to escaping local optima is also time-consuming. Thus, developing an effective UL-based solver for a single instance is crucial without relying on training data ${\mathcal{D}}$. Our numerical experiments, detailed in Appendix F.1, also validated this optimization issue, demonstrating that as the problem complexity increases, PI-GNN solver tends to be absorbed into the trivial local optima ${\bm{p}}_{{\bm{\theta}}}={\bm{0}}_{N}$ in some problems, resulting in prolonged plateaus which significantly decelerates the learning process and, in particularly challenging cases, can render learning entirely infeasible.
@@ 3.2 Continuous relaxation annealing
To address these issues, we propose a penalty term to control the discreteness and continuity of the relaxed variables, formulated as follows:
@@
Refer to Appendix B.1 for the detailed proof. Theorem 3.1 can be generalized for any convex function $\Phi({\bm{p}};C)$ which has a unique maximum at $\nicefrac{{{\bm{1}}_{N}}}{{2}}$ and becomes a global minimum for all ${\bm{p}}\in\{0,1\}^{N}$, e.g., binary cross entropy $\Phi_{\mathrm{CE}}({\bm{p}})=\sum_{i=1}^{N}(p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i}))$ introduced in Sun et al. (2022); Sanokowski et al. (2024) for the UL-based solvers using history/data. Additionally, the penalty term eliminates the stationary point ${\bm{p}}^{\ast}={\bm{0}}_{n}$ in Section 3.1, avoiding convergence to the plateau. For UL-based solvers, the penalty term is expressed as follows:
@@
Figure 1: The value $\gamma<0$ facilitates exploration by smoothing the objective function. As $\gamma$ increases, it enforces optimal discrete solutions by smoothing out suboptimal continuous ones.
On the other hand, the gradient of our penalty term in Eq. 5, $\nicefrac{{\partial\Phi({\bm{p}})}}{{\partial p_{i}}}$, is bounded within the range of $[-2\alpha,2\alpha]$ for any $\gamma$, preventing the gradient divergence issue in $\Phi_{\mathrm{CE}}({\bm{p}})$. Additionally, by increasing $\alpha$, the absolute value of the gradient near $\nicefrac{{1}}{{2}}$ becomes smaller, allowing control of the strength of smoothing out towards the discrete solution near $\nicefrac{{1}}{{2}}$. We also propose an early stopping strategy that monitors the loss function and the penalty term. The annealing and learning processes are stopped when the penalty term approaches zero, i.e., $\Phi({\bm{\theta}};C)\approx 0$. Various annealing schedules can be considered; This study employ the following scheduling strategy: $\gamma(\tau+1)\leftarrow\gamma(\tau)+\varepsilon$, where the scheduling rate $\varepsilon\in{\mathbb{R}}_{+}$ is a small constant, and $\tau$ denotes the update iterations of the trainable parameters. We refer to the PI-GNN solver with this continuous relaxation annealing as CRA-PI-GNN solver. Here, two additional hyperparameters are introduced: the initial scheduling value $\gamma(0)$ and the scheduling rate $\varepsilon$. Our numerical experiments suggest that better solutions are obtained when $\gamma(0)$ is set to a small negative value and $\varepsilon$ is kept small; the ablation study for these parameters in Appendix F.3.
@@ 4 Related Work
Previous works on UL-based solvers have addressed various problems, such as MaxCut problems (Yao et al., 2019) and traveling salesman problems (Hudson et al., 2021) relying on carefully designed problem-specific objectives. Some studies have also investigated constraint satisfaction problems (Amizadeh et al., 2018; Toenshoff et al., 2019), but applying these approaches to general CO problems typically requires problem reductions. Karalias and Loukas (2020) proposed Erdős Goes Neural (EGN) solver, unsupervised learning framework for general CO problems based on Erdős’ probabilistic method. This solver can generate solutions through inference process using training instances. Subsequently, Wang et al. (2022) proposed an entry-wise concave continuous relaxation, which generalized EGN solver to wider CO problems. On the other hand, Schuetz et al. (2022a, b) proposed PI-GNN solver, UL-based solver for a single instance, automatically learning problem-specific heuristics through training process. However, Angelini and Ricci-Tersenghi (2023); Boettcher (2023) pointed out the optimization difficulties where PI-GNN solver could not achieve results comparable to those of greedy algorithms. Wang and Li (2023) also claimed optimization issues, even for the EGN solver, and emphasized the importance of learning from training data/history to overcome local optima and then proposed Meta-EGN solvers, a meta-learning approach that updates NN network parameters for individual CO problem instances. Then, Wang and Li (2023) introduced a meta-learning method for EGN solvers, which NN parameters are updated for individual CO problem instances. To address these optimization issue, Sun et al. (2022); Sanokowski et al. (2024) proposed annealing strategy similar to simulated annealing (Kirkpatrick et al., 1983).
@@ 5 Experiments
We begin by evaluating the performance of CRA-PI-GNN solver on the MIS and the MaxCut benchmark problems across multiple graphs of varying sizes, demonstrating that CRA effectively overcomes optimization challenges without relying on data/history ${\mathcal{D}}=\{C_{\mu}\}$. We then extend the evaluation to the DBM problems, showing the applicability to more practical CO problems. For the objective functions and the detailed explanations, refer to Appendix E.1.
@@
In all experiments, the baseline methods include PI-GNN solver (Schuetz et al., 2022a) as the direct baseline of a UL-based solver for a single instance. For the MIS problems, we also consider the random greedy algorithm (RGA) and DGA (Angelini and Ricci-Tersenghi, 2019) as heuristic baselines. For the MaxCut problems, RUN-CSP solver (Toenshoff et al., 2019) is considered as an additional baseline, and a standard greedy algorithm and SDP based approximation algorithm (Goemans and Williamson, 1995) are considered as an additional classical baseline. Both are implemented by (Mehta, 2019). Note that we do not consider UL-based solvers for learning generalized heuristics (Karalias and Loukas, 2020; Wang et al., 2022; Wang and Li, 2023), which rely on training instances ${\mathcal{D}}=\{C_{\mu}\}_{\mu=1}^{P}$. The objective of this study is to evaluate whether CRA-PI-GNN solver can surpass the performance of both PI-GNN solver and greedy algorithms without any training dataset. However, for the MIS problem, EGN solver (Karalias and Loukas, 2020) and Meta-EGN solver (Wang and Li, 2023) are considered to confirm that CRA can overcome the optimization issues without training instances.
@@
The objective of the numerical experiments is to compare CRA-PI-GNN solver with PI-GNN solver. Thus, we follow the same experimental configuration described as the experiments in Schuetz et al. (2022a), employing a simple two-layer GCV and GraphSAGE (Hamilton et al., 2017) implemented by the Deep Graph Library (DGL) (Wang et al., 2019); Refer to Appendix D.1 for the detailed architectures. We use the AdamW (Kingma and Ba, 2014) optimizer with a learning rate of $\eta=10^{-4}$ and weight decay of $10^{-2}$. The GNNs are trained for up to $5\times 10^{4}$ epochs with early stopping, which monitors the loss function $\hat{l}({\bm{\theta}};{\bm{\lambda}})$ and the penalty term $\Phi({\bm{\theta}};\gamma)$ with tolerance $10^{-5}$ and patience $10^{3}$ epochs; Further details are provided in Appendix D.2. We set the initial scheduling value to $\gamma(0)=-20$ for the MIS and matching problems, and we set $\gamma(0)=-6$ for the MaxCut problems with the scheduling rate $\varepsilon=10^{-3}$ and curve rate $\alpha=2$ in Eq. (6). These values are not necessarily optimal, and refining these parameters can lead to better solutions; Refer to Appendix F.3 and Appendix F.4 for an ablation study of these parameters.
@@ 5.1 MIS problems
MIS problems on RRGs with a degree $d$ larger than $16$ is known to be hard problems (Barbier et al., 2013). As discussed in Section 3.1, Angelini and Ricci-Tersenghi (2023); Wang and Li (2023) have posted the optimization concerns on UL-based solvers. However, we call these claim into question by substantially outperforming heuristics DGA and RGA for the MIS on graphs with $d=20,100$, without training/historical instances ${\mathcal{D}}=\{G^{\mu}\}_{\mu=1}^{n}$, as shown in Table 1.
@@
We next evaluate the computational scaling of the CRA-PI-GNN solver for MIS problems with large-scale RRGs with a node degree of $100$ in Fig. 1, following previous studies (Schuetz et al., 2022a; Wang and Li, 2023). Fig. 1 demonstrated a moderate super-linear scaling of the total computational time, approximately $\sim N^{1.4}$ for GCN and $\sim N^{1.7}$ for GraphSage. This performance is nearly identical to that of the PI-GNN solver (Schuetz et al., 2022a) for problems on RRGs with lower degrees. It is important note that the runtimes of CRA-PI-GNN solver heavily depend on the optimizer for GNNs and annealing rate $\varepsilon$; thus this scaling remains largely unchanged for problems other than the MIS on $100$ RRG.

### v4

@@ 3 Continuous relaxation annealing for UL-based solvers
In this section, we discuss the practical issues associated with UL-based solvers and then introduce continuous relaxation annealing (CRA) as a proposed solution.
@@ 3.1 Motivation: practical issues of UL-based solvers
UL-based solvers (Type II) (Schuetz et al., 2022a, b) are effective in addressing large-scale CO problems. However, these solvers present following two practical issues, highlighted in several recent studies (Wang and Li, 2023; Angelini and Ricci-Tersenghi, 2023). Additionally, we numerically validate these issues; see Appendix F.1 for detailed results.
@@
UL-based solvers employ a continuous relaxation strategy to train NNs and then convert the relaxed continuous variables into discrete binary values through artificial rounding as discussed in Section 2.1. This inherent ambiguity in continuous relaxation strategy often results in potential discrepancies between the optimal solutions of the original discrete CO problem and those of the relaxed continuous one. Continuous relaxation expands the solution space, often producing continuous values that lower the cost compared to an optimal binary value. Indeed, while linear relaxation can provide an optimal solution for discrete problems on bipartite graphs (Hoffman and Kruskal, 2010), it typically results in solutions with $\nicefrac{{1}}{{2}}$ values, which is known to half-integrality (Nemhauser and Trotter Jr, 1974). Existing rounding methods (Schuetz et al., 2022b; Wang et al., 2022) often lose robustness in these scenarios. In practice, PI-GNN solver often outputs values near $\nicefrac{{1}}{{2}}$, underscoring the limitations of current rounding techniques for UL-based solvers.
@@
Recently, Angelini and Ricci-Tersenghi (2023) demonstrated that PI-GNN solver falls short of achieving results comparable to those of the degree-based greedy algorithm (DGA) (Angelini and Ricci-Tersenghi, 2019) when solving the MIS problems on RRGs. Angelini and Ricci-Tersenghi (2023) further emphasized the importance of evaluating UL-based solvers on complex CO problems, where greedy algorithms typically perform worse. A representative example is the MIS problems on RRGs with a constant degree $d>16$, where a clustering transition in the solution space creates barriers that impede optimization. Moreover, Wang and Li (2023) emphasized the importance of using training/historical datasets, ${\mathcal{D}}=\{C_{\mu}\}_{1\leq\mu\leq P}$, which contain various graphs and initialization using outputs from greedy solvers, such as DGA and RGA for MIS problems. Their numerical analysis indicated that PI-GNN solver tends to get trapped in local optima when directly optimized directly for a single instance without leveraging a training dataset ${\mathcal{D}}$. However, in a practical setting, systematic methods for generating or collecting training datasets ${\mathcal{D}}$ to effectively avoid local optima remains unclear. Additionally, training on instances that do not contribute to escaping local optima is time-consuming. Therefore, it is crucial to develop an effective UL-based solver that can operate on a single instance without relying on training data, ${\mathcal{D}}$. Our numerical experiments, detailed in Appendix F.1, also confirmed this optimization issue. They demonstrated that as problem complexity increases, the PI-GNN solver is often drawn into trivial local optima, ${\bm{p}}_{{\bm{\theta}}}={\bm{0}}_{N}$, in certain problems. This entrapment results in prolonged plateaus that significantly slow down the learning process and, in especially challenging cases, can render learning entirely infeasible. Our numerical experiments, detailed in Appendix F.1, also validated this optimization issue, demonstrating that as the problem complexity increases, PI-GNN solver tends to be absorbed into the trivial local optima ${\bm{p}}_{{\bm{\theta}}}={\bm{0}}_{N}$ in some problems, resulting in prolonged plateaus which significantly decelerates the learning process and, in particularly challenging cases, can render learning entirely infeasible.
@@ 3.2 Continuous relaxation annealing
To address these issues, we propose a penalty term to control the balance between discreteness and continuity in the relaxed variables, formulated as follows:
@@
Assuming the objective function $\hat{l}({\bm{p}};C)$ is bounded within the domain $[0,1]^{N}$, as $\gamma\to+\infty$, the relaxed solutions ${\bm{p}}^{\ast}\in\mathrm{argmin}_{{\bm{p}}}\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ converge to the original solutions ${\bm{x}}^{\ast}\in\mathrm{argmin}_{{\bm{x}}}l({\bm{x}};C,{\bm{\lambda}})$. Moreover, as $\gamma\to-\infty$, the loss function $\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ becomes convex, and the relaxed solution $\nicefrac{{{\bm{1}}_{N}}}{{2}}=\mathrm{argmin}_{{\bm{p}}}\hat{r}({\bm{p}},C,{\bm{\lambda}},\gamma)$ is unique.
For the detailed proof, refer to Appendix B.1. Theorem 3.1 can be generalized for any convex function $\Phi({\bm{p}};C)$ that has a unique maximum at $\nicefrac{{{\bm{1}}_{N}}}{{2}}$ and achieves a global minimum for all ${\bm{p}}\in\{0,1\}^{N}$; an example is binary cross entropy $\Phi_{\mathrm{CE}}({\bm{p}})=\sum_{i=1}^{N}(p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i}))$, introduced by Sun et al. (2022); Sanokowski et al. (2024) for the UL-based solvers (Type I). Additionally, the penalty term eliminates the stationary point ${\bm{p}}^{\ast}={\bm{0}}_{N}$ described in Section 3.1, preventing convergence to a plateau. For UL-based solvers, the penalty term is expressed as follows:
@@
Figure 1: Annealing strategy. When $\gamma<0$, it facilitates exploration by reducing the non-convexity of the objective function. As $\gamma$ increases, it promotes optimal discrete solutions by smoothing away suboptimal continuous ones.
We propose an annealing strategy that gradually anneals the penalty parameter $\gamma$ in Eq. (6). Initially, a negative gamma value, i.e., $\gamma<0$, is chosen to leverage the properties, facilitating broad exploration by smoothing the non-convexity of $\hat{l}({\bm{\theta}};C,{\bm{\lambda}})$ and eliminating the stationary point ${\bm{p}}^{\ast}={\bm{0}}_{N}$ to avoid the plateau, as discussed in Section 3.1. Subsequently, the penalty parameter $\gamma$ is gradually increased to a positive value, $\gamma>0$, with each update of the trainable parameters (one epoch), until the penalty term approaches zero, i.e., $\Phi({\bm{\theta}},C)\approx 0$, to automatically round the relaxed variables by smoothing out suboptimal continuous solutions oscillating between $1$ or 00. A conceptual diagram of this annealing process is shown in Fig. 1.
@@
We also propose an early stopping strategy that monitors both the loss function and the penalty term, halting the annealing and learning processes when the penalty term approaches zero, i.e., $\Phi({\bm{\theta}};C)\approx 0$. Various annealing schedules can be considered; in this study, we employ the following scheduling: $\gamma(\tau+1)\leftarrow\gamma(\tau)+\varepsilon$, where the scheduling rate $\varepsilon\in{\mathbb{R}}_{+}$ is a small constant, and $\tau$ denotes the update iterations of the trainable parameters. We refer to the PI-GNN solver with this continuous relaxation annealing as CRA-PI-GNN solver. Here, two additional hyperparameters are introduced: the initial scheduling value $\gamma(0)$ and the scheduling rate $\varepsilon$. Numerical experiments suggest that better solutions are obtained when $\gamma(0)$ is set to a small negative value and $\varepsilon$ is kept low. The ablation study are presented in Appendix F.5.
@@ 4 Related Work
Previous works on UL-based solvers have addressed various problems, such as MaxCut problems (Yao et al., 2019) and traveling salesman problems (Hudson et al., 2021), using carefully tailored problem-specific objectives. Some studies have also explored constraint satisfaction problems (Amizadeh et al., 2018; Toenshoff et al., 2019), but applying these approaches to broader CO problems often requires problem-specific reductions. Karalias and Loukas (2020) proposed Erdős Goes Neural (EGN) solver, an UL-based solver for general CO problems based on Erdős’ probabilistic method. This solver generate solutions through an inference process using training instances. Subsequently, Wang et al. (2022) proposed an entry-wise concave continuous relaxation, broadening the EGN solver to a wide range of CO problems. In contrast, Schuetz et al. (2022a, b) proposed PI-GNN solver, an UL-based solver for a single CO problems that automatically learns problem-specific heuristics during the training process. However, Angelini and Ricci-Tersenghi (2023); Boettcher (2023) pointed out the optimization difficulties where PI-GNN solver failed to achieve results comparable to those of greedy algorithms. Wang and Li (2023) also claimed optimization issues with PI-GNN solver, emphasizing the importance of learning from training data and history to overcome local optima. They then proposed Meta-EGN solvers, a meta-learning approach that updates NN network parameters for individual CO problem instances. Furthermore, to address these optimization issue, Lin et al. (2023); Sun et al. (2022); Sanokowski et al. (2024) proposed annealing strategy similar to simulated annealing (Kirkpatrick et al., 1983).
@@ 5 Experiments
We begin by evaluating the performance of CRA-PI-GNN solver on the MIS and the MaxCut benchmark problems across multiple graphs of varying sizes, demonstrating that CRA effectively overcomes optimization challenges without relying on data/history ${\mathcal{D}}$. We then extend the evaluation to the DBM problems, showing the applicability to more practical CO problems. For the objective functions and the detailed explanations, refer to Appendix E.1.
5.1 Experimental settings
@@
In all experiments, the baseline methods include the PI-GNN solver (Schuetz et al., 2022a) as the direct baseline of a UL-based solver for a single instance. For the MIS problems, we also consider the random greedy algorithm (RGA) and DGA (Angelini and Ricci-Tersenghi, 2019) as heuristic baselines. For the MaxCut problems, RUN-CSP solver (Toenshoff et al., 2019) is considered as an additional baseline, and a standard greedy algorithm and SDP based approximation algorithm (Goemans and Williamson, 1995) are considered as an additional classical baseline. The parameters for the Goemans-Williamson (GW) approximation are all set according to the settings in Schuetz et al. (2022b). The implementation used the open-source CVXOPT solver with CVXPY (Mehta, 2019) as the modeling interface. Note that we do not consider UL-based solvers for learning generalized heuristics (Karalias and Loukas, 2020; Wang et al., 2022; Wang and Li, 2023), which rely on training instances ${\mathcal{D}}=\{C_{\mu}\}_{\mu=1}^{P}$. The primary objective of this study is to evaluate whether CRA-PI-GNN solver can surpass the performance of both PI-GNN solver and greedy algorithms. However, for the MIS problem, EGN solver (Karalias and Loukas, 2020) and Meta-EGN solver (Wang and Li, 2023) are considered to confirm that CRA can overcome the optimization issues without training instances.
@@
The objective of the numerical experiments is to compare the CRA-PI-GNN solver with the PI-GNN solver. Thus, we follow the same experimental configuration described as the experiments in Schuetz et al. (2022a), employing a simple two-layer GCV and GraphSAGE (Hamilton et al., 2017) implemented by the Deep Graph Library (Wang et al., 2019); Refer to Appendix D.1 for the detailed architectures.
We use the AdamW (Kingma and Ba, 2014) optimizer with a learning rate of $\eta=10^{-4}$ and weight decay of $10^{-2}$. The GNNs are trained for up to $5\times 10^{4}$ epochs with early stopping, which monitors the summarized loss function $\sum_{s=1}^{S}\hat{l}(P_{:,s})$ and the entropy term $\Phi(P;\gamma,\alpha)$ with tolerance $10^{-5}$ and patience $10^{3}$ epochs; Further details are provided in Appendix D.2. We set the initial scheduling value to $\gamma(0)=-20$ for the MIS and matching problems, and we set $\gamma(0)=-6$ for the MaxCut problems with the scheduling rate $\varepsilon=10^{-3}$ and curve rate $\alpha=2$ in Eq. (6). These values are not necessarily optimal, and refining these parameters can lead to better solutions; Refer to Appendix F.5 and Appendix F.6 for an ablation study of these parameters. Once the training process is complete, we apply projection heuristics to map the obtained soft solutions back to discrete solutions using simple projection, where for all $i\in[N]$, we map $p_{\theta,i}$ into 00 if $p_{\theta,i}\leq 0.5$ and $p_{\theta,i}$ into $1$ if $p_{\theta,i}>0.5$. However, due to the early stopping in Section 3.2, the CRA-PI-GNN solver ensures that for all benchmark CO problems, the soft solution at the end of the training process became 0 or 1 within the 32-bit Floating Point range in Pytorch GPU; thus, it is robust against a given threshold, which we set to $0.5$ in our experiments. Additionally, no violations of the constraints were observed in our numerical experiments. Thus, following results presented in are feasible solutions.
@@ 5.2 MIS problems
MIS problems on RRGs with a degree $d$ larger than $16$ is known to be hard problems (Barbier et al., 2013). As discussed in Section 3.1, Angelini and Ricci-Tersenghi (2023); Wang and Li (2023) have posted the optimization concerns on UL-based solvers. However, we call these claim into question by substantially outperforming heuristics DGA and RGA for the MIS on graphs with $d=20,100$, without training/historical instances ${\mathcal{D}}=\{G^{\mu}\}_{\mu=1}^{n}$, as shown in Table 1. See Appendix 6 for the results of solving all other Gsets, where consistently, CRA-PI-GNN provides better results as well. A comparison of the sampling-based solvers, RL-based solvers, SL-based solvers, Gurobi, and MIS-specific solvers is presented in Appendix F.2.
@@
We next evaluate the computational scaling of the CRA-PI-GNN solver for MIS problems with large-scale RRGs with a node degree of $100$ in Fig. 1, following previous studies (Schuetz et al., 2022a; Wang and Li, 2023). Fig. 1 demonstrated a moderate super-linear scaling of the total computational time, approximately $\sim N^{1.4}$ for GCN and $\sim N^{1.7}$ for GraphSage. This performance is nearly identical to that of the PI-GNN solver (Schuetz et al., 2022a) for problems on RRGs with lower degrees. It is important note that the runtimes of CRA-PI-GNN solver heavily depend on the optimizer for GNNs and annealing rate $\varepsilon$; thus this scaling remains largely unchanged for problems other than the MIS on $100$ RRG. Additionally, CRA demonstrate that the runtime remains nearly constant as graph order and density increase, indicating effective scalability with denser graphs which is presented in Appendix F.2.